OpenStudy (anonymous):

find prime factorization of 22! ?

OpenStudy (anonymous):

1x2x11

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=factorize+22%21

OpenStudy (anonymous):

no it is factorial of 22 which is needed

OpenStudy (anonymous):

nvm

OpenStudy (hba):

Use a calculator @fhmm

OpenStudy (anonymous):

first of all list the prime numbers less than or equal to 22 and then one should find the index of each such prime numbers

OpenStudy (anonymous):

for eg to find the highest power of 2 tha divides 22! we have [22/2] +[22/4]+[22/8]+[22/16]+[22/32] =11+5+2+1=19 { [ ] denotes greatest integer function} thus 2^19 divides 22! similarly one can find out for 3,5,7,11,13,17,19